Exercise 2:


1- Two trains will travel on the same straight rail and head towards each other. Speed ​​of one of them is 130Km / h and the speed of the second is 0K / h 10 for the rail. About 3km away, the two drivers both noticed the error, and each began to halt their train. If each train slowed by 1 ° m / s, would they collide? 2- At a distance d in front of him, a driver of a train with speed V sees another train


Ethe V2 and head


It runs on the same straight rail in the same direction as the first train. The driver is running


The brakes


The train is slowing down by a constant a. What is the condition that the two trains do not collide?


1
Expert's answer
2021-02-22T10:22:17-0500

"u=v_1+v_2\\t=\frac{u}{a}=\frac{130+100}{3.6\cdot1}=36.9\ s"

"s=ut-0.5at^2\\s=\frac{130+100}{3.6}(36.9)-0.5(1)(36.9)^2=2041\ m"

It is less than 3 km. Thus, the trains will not collide.